# A Local Projection Stabilised HHO Method for the Oseen Problem

### v1

@@ 6 Conclusions
In this article, we have considered a local projection stabilised Hybrid High-Order method for the Oseen problem. We have proved a strong stability result for our discrete formulation under the SUPG norm. The error analysis shows the optimal order of convergence $(k+1/2)$, assuming that the velocity and the pressure have regularity $H^{k+1}$ globally and $H^{k+2}$ locally. This analysis can be extended to three dimensions with an appropriate change in the mesh structure. The first numerical experiment shows that the optimal order of convergence is achieved. The second example shows that the boundary layer is also captured. It is our presumption that the issue of suboptimal convergence rate in example 5.2 with $\epsilon\leq 10^{-3}$ can be resolved using an adaptive approach. This will be considered in future work.
@@ Acknowledgements
The first author acknowledges the DST-SERB-MATRICS grant MTR/2023/000681 for financial support.
@@ 7 Appendix
$\sum_{T\in\mathcal{T}_{h}}(D_{T}^{k}\underline{\boldsymbol{v}}_{T},1)_{T}=\sum_{T\in\mathcal{T}_{h}}\sum_{F\in\mathcal{F}_{T}}(\boldsymbol{v}_{F}\cdot\boldsymbol{n},1)_{F}=0.$
Along with this $D_{h}^{k}$ is also a continuous linear surjective operator from $\underline{\boldsymbol{U}}_{h,0}^{k}$ to $P_{h}^{k}$. Moreover, we have the following lemma.
@@
There exists $\alpha>0$ independent of meshsize such that for all $p_{h}\in P_{h}^{k}$ we have the following
$\sup_{\underline{\boldsymbol{v}}_{h}\in\underline{\boldsymbol{U}}_{0,h}^{k}}\frac{B_{h}(\underline{\boldsymbol{v}}_{h},p_{h})}{\left\|\underline{\boldsymbol{v}}_{h}\right\|_{{\rm dG}}+(\sum_{T\in\mathcal{T}_{h}}\left\|\underline{\boldsymbol{v}}_{T}\right\|^{2}_{T})^{1/2}}\geq\alpha\left\|p_{h}\right\|.$ (7.1)
@@
It is well known that the velocity/pressure pair $[H^{1}_{0}(\Omega)]^{2}/L^{2}_{0}(\Omega)$ is stable and the continuous divergence operator is surjective. Therefore, for any $p_{h}\in P_{h}^{k}$ one can have a unique $\boldsymbol{v}\in\boldsymbol{V}$ such that $\left\|\boldsymbol{v}\right\|_{1}\leq C\left\|p_{h}\right\|$ and ${\rm div~{}}\boldsymbol{v}=-p_{h}$. Using this and the relation $\pi_{T}^{k}({\rm div~{}}\boldsymbol{v})=D_{T}^{k}(\underline{\boldsymbol{I}}_{h}(\boldsymbol{v}))$ we get
@@
$\left\|\underline{\boldsymbol{I}}_{h}(\boldsymbol{v})\right\|^{2}_{{\rm dG}}+\sum_{T\in\mathcal{T}_{h}}\left\|\underline{\boldsymbol{I}}_{T}(\boldsymbol{v})\right\|^{2}_{T}$
$=\sum_{T\in\mathcal{T}_{h}}\Big{(}\left\|\underline{\boldsymbol{I}}_{T}(\boldsymbol{v})\right\|^{2}_{T}+\left\|\nabla\underline{\boldsymbol{I}}_{T}(\boldsymbol{v})\right\|^{2}_{T}+\sum_{F\in\mathcal{F}_{T}}\int_{F}\frac{1}{h_{F}}(\pi_{F}^{k}(\boldsymbol{v})-\pi_{T}^{k}(\boldsymbol{v}))^{2}\Big{)}$
$\leq C\left\|\boldsymbol{v}\right\|_{1}^{2}\leq C\left\|p_{h}\right\|^{2}$ (7.3)
@@
Since $P_{h}^{k}$ is a reflexive space and the discrete divergence operator $D_{h}$ is surjective, the converse of the Lemma A.42 in Ern:2004:FEMbook says that for any $p_{h}\in P_{h}^{k}$ there exists $\underline{\boldsymbol{v}}_{h}\in\underline{\boldsymbol{U}}_{0,h}^{k}$ such that
$D_{h}^{k}(\underline{\boldsymbol{v}}_{h})=p_{h}\quad\text{and}\quad\left\|\underline{\boldsymbol{v}}_{h}\right\|_{{\rm dG}}^{2}+\sum_{T\in\mathcal{T}_{h}}\left\|\boldsymbol{v}\right\|^{2}_{T}\leq C\left\|p_{h}\right\|^{2},~{}\text{C depends on }\Omega.$ (7.4)

### v2

@@ 6 Conclusions
In this article, we have considered a local projection stabilised Hybrid High-Order method for the Oseen problem. We have proved a strong stability result for our discrete formulation under the SUPG norm. The error analysis shows the optimal order of convergence $(k+1/2)$, assuming that velocity and pressure have regularity $H^{k+1}$ globally and $H^{k+2}$ locally. In the second example with very small $\epsilon$, we obtain suboptimal convergence rates for both the HHO-LPS method and the HHO method discussed in [1]. It is our presumption that the issue of the suboptimal convergence rate in example 5.2 with very small $\epsilon$ can be resolved using an adaptive approach. This will be considered in future work.
@@ Acknowledgements
The first author acknowledges the DST-SERB-MATRICS grant MTR/2023/000681 for financial support and DST FIST grant SR/FST/MS II/2023/139 for partial financial support.
@@ 7 Appendix
$\sum_{T\in\mathcal{T}_{h}}(D_{T}^{k}\underline{\boldsymbol{v}}_{T},1)_{T}=\sum_{T\in\mathcal{T}_{h}}\sum_{F\in\mathcal{F}_{T}}(\boldsymbol{v}_{F}\cdot\boldsymbol{n}_{TF},1)_{F}=0.$
Along with this, $D_{h}^{k}$ is also a continuous linear surjective operator from $\underline{\boldsymbol{U}}_{h,0}^{k}$ to $P_{h}^{k}$. Moreover, we have the following lemma.
@@
There exists $\alpha>0$ independent of the meshsize such that for all $p_{h}\in P_{h}^{k}$ we have the following
$\sup_{\underline{\boldsymbol{v}}_{h}\in\underline{\boldsymbol{U}}_{0,h}^{k}}\frac{B_{h}(\underline{\boldsymbol{v}}_{h},p_{h})}{\left\|\underline{\boldsymbol{v}}_{h}\right\|_{1,h}}\geq\alpha\left\|p_{h}\right\|.$ (7.1)
@@
It is well known that the velocity/pressure pair $[H^{1}_{0}(\Omega)]^{d}/L^{2}_{0}(\Omega)$ is stable and the continuous divergence operator is surjective. Therefore, for any $p_{h}\in P_{h}^{k}$ one can have a unique $\boldsymbol{v}\in\boldsymbol{V}$ such that $\left\|\boldsymbol{v}\right\|_{1}\leq C\left\|p_{h}\right\|$ and ${\rm div~{}}\boldsymbol{v}=-p_{h}$. Using this and the relation $\pi_{T}^{k}({\rm div~{}}\boldsymbol{v})=D_{T}^{k}(\underline{\boldsymbol{I}}_{h}(\boldsymbol{v}))$ we get
@@
$\left\|\underline{\boldsymbol{I}}_{h}(\boldsymbol{v})\right\|^{2}_{1,h}$ $=\sum_{T\in\mathcal{T}_{h}}\Big{(}\left\|\nabla\underline{\boldsymbol{I}}_{T}(\boldsymbol{v})\right\|^{2}_{T}+\sum_{F\in\mathcal{F}_{T}}\int_{F}\frac{1}{h_{F}}(\pi_{F}^{k}(\boldsymbol{v})-\pi_{T}^{k}(\boldsymbol{v}))^{2}\Big{)}$
$\leq C\left\|\boldsymbol{v}\right\|_{1}^{2}\leq C\left\|p_{h}\right\|^{2}.$ (7.3)
@@
Since $P_{h}^{k}$ is a reflexive space and the discrete divergence operator $D_{h}$ is surjective, the converse of the Lemma A.42 in [40] says that for any $p_{h}\in P_{h}^{k}$ there exists $\underline{\boldsymbol{v}}_{h}\in\underline{\boldsymbol{U}}_{0,h}^{k}$ such that
$D_{h}^{k}(\underline{\boldsymbol{v}}_{h})=p_{h}\quad\text{and}\quad\left\|\underline{\boldsymbol{v}}_{h}\right\|_{1,h}^{2}\leq C\left\|p_{h}\right\|^{2},~{}\text{C depends on }\Omega.$ (7.4)